OpenStudy (anonymous):

determine weather the sum of (3sin(n))/n^(5/3) is convergent or divergent from 0 to infinity

OpenStudy (anonymous):

i would imagine it converges

OpenStudy (anonymous):

since n^a, a is bigger than 1

OpenStudy (anonymous):

it is convergent

OpenStudy (anonymous):

well \[\frac{5}{3}>1\][

OpenStudy (anonymous):

and so \[\sum\frac{1}{n^{\frac{5}{3}}}\] converges

hero (hero):

I don't reallly know much about converge and diverge. I just found series expansion on wolfram alpha. I see that it is not reliable

OpenStudy (anonymous):

and \[\sin(n)\leq 1\]

OpenStudy (anonymous):

wow i can't even read that. but we can use wolfram to check

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sum++sin%28n%29%2Fn^%283%2F5%29 yeah not only does it converge, but it is less than two

hero (hero):

So that part that is posted on wolfram alpha that I posted, what do you call that?

OpenStudy (anonymous):

so what is the first step? can i do the comparison test?

hero (hero):

I will study these more so that I know the steps.

OpenStudy (anonymous):

yes you can compare it to a "p" series ( i think that is what it is called)

OpenStudy (anonymous):

\[\sum\frac{1}{n^p}\] converges if \[p>1\]

OpenStudy (amistre64):

\[ \frac{3sin(n)}{n^{5/3}}\] sin(n) is alternating

OpenStudy (anonymous):

well not exactly

OpenStudy (amistre64):

almost got that right lol

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

cos(n) alternated ;)

OpenStudy (amistre64):

...pi*n that is , ill get my head on straight eventually

OpenStudy (anonymous):

but in any case it is true that \[\sin(n)\leq 1\]so \[\frac{\sin(n)}{n^{\frac{5}{3}}}\leq \frac{1}{n^{\frac{5}{3}}}\]

OpenStudy (anonymous):

one on the right converges so one on the left does as well.

OpenStudy (anonymous):

yeah that pi ...